A moving sphere has a head-on elastic collision with an initially stationary sphere. After collision, the kinetic energies of the two spheres are equal. Show that the mass ratio of the two spheres is 0.1716?
1
Expert's answer
2020-06-18T10:59:53-0400

As far as the collision is elastic and there are no external forces on the system, the conservation laws of momentum and kineticenergy:


"m_1 v_1 = m_1v_1' + m_2 v_2'\\
\dfrac{m_1v_1^2}{2} = \dfrac{m_1v_1'^2}{2} + \dfrac{m_2v_2'^2}{2}"

where "v_1" is the speed of the first sphere before the collision and "v_1'" and "v_2'" are the speed of the first and the second spheres respectively. As far, as the kinetic energies of the two spheres are equal after collision, we can write:


"\dfrac{m_1v_1'^2}{2} = \dfrac{m_2v_2'^2}{2}\\"

Substituting this into the second equation, get:


"\dfrac{m_1v_1^2}{2} = \dfrac{m_1v_1'^2}{2} + \dfrac{m_2v_2'^2}{2} =\dfrac{m_1v_1'^2}{2} + \dfrac{m_1v_1'^2}{2} = m_1v_1'^2"

Thus:

"\dfrac{m_1v_1^2}{2} = m_1v_1'^2\\
v_1' = \dfrac{v_1}{\sqrt{2}}"

Let's substitute this to the first equation (momentum conservation) and express "v_2'":


"m_1 v_1 = m_1\dfrac{v_1}{\sqrt{2}}+ m_2 v_2'\\"

"v_2' = \dfrac{m_1v_1(\sqrt{2}-1)}{\sqrt{2}m_2}"

Again, from the eqation "\dfrac{m_1v_1'^2}{2} = \dfrac{m_2v_2'^2}{2}" note, that:

"\dfrac{m_2}{m_1} = \dfrac{v_1'^2}{v_2'^2}"

Substituting here the expressions for "v_1'" and "v_2'" obtained before, get:


"\dfrac{m_2}{m_1} = \dfrac{v_1^2}{2}\cdot \dfrac{2m_2^2 }{m_1^2v_1^2(\sqrt{2}-1)^2}"

After expressing the ratio "m_2\/m_1", get:


"\dfrac{m_2}{m_1} = (\sqrt{2}-1)^2 \approx 0.1716"

QED.